The ratio of boys to girls in a school is 2:3. If the school has 632 students, about how many are girls?
Answer by TimothyLamb(4379)

b/g = 2/3
---
b = (2/3)g
b + g = 632
---
b + g = 632
(2/3)g + g = 632
(2/3)g + (3/3)g = 632
(5/3)g = 632
g = 632*(3/5)
g = 379.2
---
answer:
about 379 girls
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---